OpenStudy (anonymous):

plz help simplify 10/x-5 divided by 5

OpenStudy (anonymous):

@Paynesdad

OpenStudy (anonymous):

my answe choices are a.2-x/x b.1/x c.1 d.2

OpenStudy (abb0t):

@BulletWithButterflyWings

OpenStudy (anonymous):

i am usually good at this kind of stuff but i am stuck on how to do this problem

OpenStudy (mathstudent55):

Is this the problem? \( \dfrac{10}{x - 5} \div 5 \)

OpenStudy (anonymous):

yes it is

OpenStudy (mathstudent55):

To divide, you multiply by the reciprocal: \( =\dfrac{1}{x - 5} \times \dfrac{1}{5} \) \( =\dfrac{1}{5(x - 5) } \)

OpenStudy (anonymous):

ok i see how you did that

OpenStudy (anonymous):

so would my answer be 2-x/x

OpenStudy (anonymous):

is my answer right?????

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

I gave you the answer already, why are you asking for the answer?

OpenStudy (mathstudent55):

2 - x/x has nothing to do with this problem.

OpenStudy (anonymous):

but the answer you gave me isnt one of my choices

OpenStudy (anonymous):

my answe choices are a.2-x/x b.1/x c.1 d.2

OpenStudy (anonymous):

can you help me @robtobey

OpenStudy (anonymous):

\[\frac{1}{5} \left(\frac{10}{x}-5\right)\text{=}\frac{2-x}{x}\]\[\left(\frac{1}{5}\right)\frac{10}{x-5}=\frac{2}{x-5} \]

OpenStudy (anonymous):

ok thanks for your help

OpenStudy (anonymous):

You're welcome